# Enhancing Brain Source Reconstruction through Physics-Informed 3D Neural Networks

### v1

@@ Abstract
Index Terms:
Deep Learning, EEG, Inverse Problem, Neuroimaging, Source Localisation
@@ I Introduction
Figure 1: Overview of different approaches for source localization from EEG data. Top The classical (Minimum Norm) approaches uses the pseudo-inverse ${\textbf{L}^{\dagger}}$ of the forward model to estimate the sources $\hat{\textbf{x}}$ from the EEG measurements y. Middle: Data-driven end-to-end approaches train some neural network $g_{\theta}$ on simulated data, where the source activity x is drawn from the data prior $p(\textbf{x})$. During inference, the network predicts the sources $\hat{\textbf{x}}$ directly from measurements without explicitly using the forward model. Bottom: Our approach, 3D-PIUNet, integrates data-driven learning with physics-informed modeling by incorporating the forward model’s physical information through the pseudo-inverse. After transforming the EEG measurements into a grid-like source space, a 3D convolutional U-Net $f_{\theta}$ refines the pseudo-inverse solution according to the learned data prior, resulting in an improved source estimate $\hat{x}$.
@@
To understand both normal brain function and dysfunction, it is essential to accurately identify the precise sources of electrical activity within the brain, a process known as source localization[1]. This is an important tool in cognitive neuroscience, enabling the study of brain responses to various stimuli and the functional organization of the brain. Electroencephalography (EEG) and magnetoencephalography (MEG) are non-invasive techniques that allow for monitoring brain activity without the need for surgical intervention [2]. Moreover, M/EEG offers high temporal resolution, capturing neural events on the order of milliseconds, which is critical for studying fast neural dynamics and the timing of cognitive processes. This high temporal resolution makes EEG particularly suitable for investigating the rapid neural processes underlying cognition, perception, and action.
@@
However, pinpointing the precise brain regions responsible for these signals remains a challenge due to the ill-posed nature of the EEG inverse problem [3]. Conventional methods try to address this by posing priors on the source activation, like sparsity or minimal norm activity [4]. Manual crafted priors bias the solution to specific settings. For example sparse methods do not work on distributed sources and minimum norm methods do not work for sparse sources. In recent years, neural networks [5, 6, 7, 8] have shown improved performance by learning a direct mapping from the measurement to sources in an end-to-end fashion. Unlike traditional techniques, neural networks offer greater flexibility in formulating priors on brain activation by simulated data, that serves as a proxy for real-world conditions, possibly capturing a wide range of brain activity patterns. However, these approaches typically incorporate the forward model information only indirectly through their training data, which can limit their accuracy and generalizability in different contexts [9].
@@
To combine worlds, we propose a novel hybrid approach, named 3D-PIUNet, that combines the strengths of both classical and deep learning methods. We incorporate physics-informed prior knowledge from the forward model via the pseudo-inverse solution provided by eLORETA [10]. Our method leverages a 3D convolutional neural network to refine this pseudo-inverse solution according to a learned data-driven prior that is able to capture complex patterns in the source space. We show a comparison of classical approaches, end-to-end learning and our proposed 3D-PIUNet in Fig. 1.
@@
Our experimental results on a comprehensive synthetic dataset demonstrate that 3D-PIUNet consistently outperforms both eLORETA and end-to-end deep learning methods in terms of source localization accuracy. The network effectively learns to refine the pseudo-inverse solution, leading to more accurate and robust reconstructions, especially in the presence of noise. This work highlights the potential of best-of-both-worlds approaches to overcome the limitations of traditional and data-driven end-to-end methods in EEG source imaging.
@@
  • extensively benchmark our approach on diverse settings, showing superiority to both classical methods and end-to-end approaches across different signal-to-noise levels, number and size of sources.
@@ II-A Problem Setup
where a known lead field matrix $\textbf{L}\in\mathbb{R}^{M\times N}$ maps the electrical activity of the brain, $\textbf{x}\in\mathbb{R}^{N\times T}$, to the sensor measurements, $\textbf{y}\in\mathbb{R}^{M\times T}$. Here, $M$ describes the number of sensors on the scalp, $N$ the number of potential brain sources within the brain and $T$ the number of time steps. The noise vector $\boldsymbol{\epsilon}$ describes the measurement noise, covering inaccuracies in the forward model, electrical pollution and sensor errors.
@@ II-B Classical Approaches to EEG Source Localization
On the other hand, distributed methods assume that brain activity can be spread across regions involving multiple dipoles. These methods place a dense set of potential source locations within the head volume. The inverse problem becomes a regularized regression problem, predicting the amplitudes of each location, thus not constraining the number of dipoles [14]. The Minimum Norm Estimates (MNE) [4] estimates the distribution of sources by minimizing the $\ell_{2}$-norm of the source amplitudes. The $\ell_{2}$-norm offers a closed-form solution and is, therefore, fast to compute. However, it has a bias towards sources close to the sensors. To counteract this bias, different weighted minimum norm solutions [15] have been proposed, such as dSPM [16], LORETA [17], sLORETA [18] and eLORETA [10] or combinations of those [19]. As this weighting promotes a spatial smoothness prior, small sources might be missed or blurred. To obtain sparse solution with minimal non-zero sources, one can replace the $\ell_{2}$-norm with $\ell_{1}$-norm, called Least Absolute Shrinkage and Selection Operator (LASSO) [20] or selective minimum norm [21]. By combination of sparse and dense prior, it is also possible to generate solutions with mixed spatial extent [22].
@@ II-C Deep Learning Methods
In recent years, learning-based approaches have become popular in the sciences due to their flexibility and fast inference time, including EEG data analysis [29]. A better understanding of brain activity due to virtual brain models [30] and large-scale dynamical models [31] allow the generation of realistic simulation data acting as a powerful data-prior. To learn this data prior, previous work focused on end-to-end learning with different network architectures. For single timestep predictions, fully connected networks have been used to predict source locations directly [32, 33] showing the same limitation in fixed number of sources as dipole fitting. Convolutional networks have been proposed to encode spatial correlation of sensor measurement [7], mixing EEG and MEG signals [34] or to incorporate the time dimension [35]. Other recent approaches use LSTMs for temporal predictions after encoding the measurements either with residual networks [6], graph convolutional networks [36], dilated convolution [37] or multi-head attention [38]. However, all these end-to-end methods use the forward model only during data generation and require retraining or fine-tuning when the forward model changes [9]. Physics-informed machine learning bridges the gap and has shown superior performances in fields like fluid dynamics [39], quantum chemistry [40], quantum circuits [41], and other inverse problems like Computed Tomography (CT) [42] and MRI [43]. In the following, we explain our hybrid approach 3D-PIUNet for source localization, combining the benefits of classical and deep learning methods.
@@ III-A Starting from Pseudo-Inverse Solution
Previous works using neural networks for source estimation [7, 6, 38] learn a mapping $\hat{\textbf{x}}=g_{\theta}(\textbf{y}):\mathbb{R}^{M}\to\mathbb{R}^{N}$ from sensor to source space. Even though this end-to-end learning is common practice in modern deep learning, it has some drawbacks for inverse problems. Most importantly, the model needs to learn the physical dynamics only from data, while not having access to a forward model. Therefore, the learned model heavily depends on the lead fields used during training and cannot generalize well to other forward models. For example, when the number of sensors change, classical end-to-end networks require the training of a new input layer.
We tackle these shortcomings by starting from a pseudo-inverse solution that maps the measurements to the source space. This way, the network can leverage the established physical knowledge embedded in the forward model. Importantly, as long as the pseudo-inverse matrix maintains full column rank, this transformation ensures that no information is lost. During training, the model learns a data prior in source space and can focus on refining the pseudo-inverse solution according to this prior. While this framework has been applied to other inverse problems like CT [42] and MRI [43] reconstruction, to the best of our knowledge, it is the first application to EEG data. Given the pseudo-inverse ${\textbf{L}^{\dagger}}$, the neural network is trained using the transformed measurements $\tilde{\textbf{x}}={\textbf{L}^{\dagger}}\textbf{y}$ as an input. It then learns the refinement task as a mapping within source space $f_{\theta}:\mathbb{R}^{N}\to\mathbb{R}^{N}$:
@@
One can interpret the pseudo-inverse solution as a non-learnable physics-informed input layer that precedes the deep neural network. This way, the pseudo-inverse can be adapted to account for different subjects at inference time. By mapping the measurements into the source space, the input dimensionality does not depend on the sensing system and the same architecture can be used, enabling transfer learning.
@@
We divide the brain sources to a regular 3D grid. This allows the use of a 3D convolutional neural network (CNN) [45]. CNNs excel in this domain due to their parameter efficiency and ability to capture spatial relationships. For tasks like source localization, where preserving spatial information across scales is crucial, the U-Net architecture [46] is particularly well-suited. It employs residual connections between the encoding (downscaling) and decoding (upscaling) blocks, allowing information from the input to propagate directly to the decoder and aiding in the reconstruction of fine details. For this work, we adopt a state-of-the-art U-Net architecture used in denoising diffusion models [47].
@@
Before applying the network to our pseudo sources $\tilde{x}$, we use zero padding, extending the potential source locations to a regular $32\times 32\times 32$ grid. The architecture begins by lifting the pseudo inverse to a $C=32$ channel feature map, effectively creating multiple representations of the initial source estimate. Next, the network progressively down-samples to a bottleneck resolution of ($8\times 8\times 8$) while doubling the feature channels after each down convolution, increasing the receptive field of each convolutional layer. At each resolution, we use two 3D convolutional residual blocks. Each residual block starts with a Group Norm layer [48], a normalization technique that is less dependent on batch size, leading to improved training stability and performance. This is followed by SiLU activation, a smooth, non-linear activation function outperforming ReLU for deep neural networks [49]. The core computation is done by the $3\times 3\times 3$ convolution layers that aggregates neighboring information to learn descriptive features and refine the source estimates. Each block uses residual connections, where the input of each block is added to its output. This improves the gradient flow during backpropagation and simplifies the optimization problem [50].
@@
At the bottleneck, we employ attention layers [51]. In contrast to Vision Transformers [52] that represent a group of pixels as a single token, we use a token for each voxel, but in low-resolution embedding space. The attention mechanism allows the model to focus on the most relevant regions for source reconstruction. These lower layers of the network capture a broader spatial context, and attention allows the model to weigh the importance of different regions within this context, leading to more accurate refinement of the source estimates.
@@ III-C Training Details
$\text{Loss}(\textbf{x},f_{\theta}({\textbf{L}^{\dagger}}\textbf{y}))=||\textbf{x}-f_{\theta}({\textbf{L}^{\dagger}}\textbf{y})||_{1}$ (3)
@@
We used the Adam optimizer [54] with a learning rate of 0.00001 and a batch size of 32. The training process consisted of 60000 steps, where the data is generated via the simulation explained in section IV-B.
@@ IV-A Forward Model
To compute the forward model, we employ a three-shell Boundary Element Method (BEM) [55] approach implemented within the MNE-Python library [56]. To represent potential sources of brain activity, a volumetric source space is constructed. This source space consists of a three-dimensional grid with a uniform voxel spacing of 7 mm. The fsaverage [57] head geometry is used with three concentric layers representing the scalp, skull, and brain tissues and standard conductivity values assigned to these layers: $0.3S/m^{2}$ (scalp), $0.006S/m^{2}$ (skull), and $0.3S/m^{2}$ (brain). These values are based on established studies that provide realistic approximations for human head tissues [7].
To calculate the forward solution, the minimum distance between potential sources and the inner skull surface is set to 5 mm. This parameter helps exclude source locations that are unrealistically close to the skull boundary, resulting in $N=4819$ sources. The volumetric source space is constructed with freely oriented sources, allowing each voxel to represent neural activity with a three-dimensional vector for direction and amplitude, providing greater flexibility in modeling complex neural activity patterns. Sensor positions are defined using a standard Easycap-M10 montage with $M=61$ sensors. The resulting lead field matrix $\textbf{L}\in\mathbb{R}^{(61\times(4819\cdot 3))}$ describes the mapping from source locations and orientation to sensors.
As using the same model for both forward and inverse solutions can lead to over-optimistic results — the so-called inverse crime [58, 59] — we generate a second BEM solution for testing. Here, the conductivity is changed to $0.332S/m^{2}$ (scalp), $0.0113S/m^{2}$ (skull), and $0.332S/m^{2}$ (brain) following an established setup [7, 60]. This ensures that our models are validated against a different forward model, simulating more realistic conditions and enhancing the robustness of our validation process.

### v2

# Enhancing Brain Source Reconstruction by Initializing 3D Neural Networks with Physical Inverse Solutions

@@ I Introduction
Figure 1: Overview of different approaches for source localization from EEG data. Top The classical (Minimum Norm) approaches uses the pseudo-inverse $\textbf{L}^{\dagger}$} of the forward model, combined with an identity mapping $\mathbf{I}$ for illustrative purposes, to estimate the sources $\hat{\textbf{x}}$ from the EEG measurements y. Middle: Data-driven end-to-end approaches train some neural network $g_{\theta}$ on simulated data, where the source activity x is drawn from the data prior $p(\textbf{x})$. During inference, the network predicts the sources $\hat{\textbf{x}}$ directly from measurements without explicitly using the forward model. Bottom: Our approach, 3D-PIUNet, integrates data-driven learning with physics-informed modeling by incorporating the forward model’s physical information through the pseudo-inverse. After transforming the EEG measurements into a grid-like source space, a 3D convolutional U-Net $f_{\theta}$ refines the pseudo-inverse solution according to the learned data prior, resulting in an improved source estimate $\hat{x}$.
@@
To understand both normal brain function and dysfunction, it is essential to accurately identify the precise sources of electrical activity within the brain, a process known as source localization[50]. This is an important tool in cognitive neuroscience, enabling the study of brain responses to various stimuli and the functional organization of the brain. Electroencephalography (EEG) and magnetoencephalography (MEG) are non-invasive techniques that allow for monitoring brain activity without the need for surgical intervention [33]. Moreover, M/EEG offers high temporal resolution, capturing neural events on the order of milliseconds, which is critical for studying fast neural dynamics and the timing of cognitive processes. This high temporal resolution makes EEG particularly suitable for investigating the rapid neural processes underlying cognition, perception, and action[22, 4].
@@
However, pinpointing the precise brain regions responsible for these signals remains a challenge due to the ill-posed nature of the EEG inverse problem [2]. Conventional methods try to address this by posing priors on the source activation, like sparsity or minimal norm activity [28]. Manual crafted priors bias the solution to specific settings. For example, sparse methods do not work on distributed sources and minimum norm methods do not work for sparse sources [14]. In recent years, neural networks [12, 67, 35, 70] have shown improved performance by learning a direct mapping from the measurement to sources in an end-to-end fashion. Unlike traditional techniques, neural networks offer greater flexibility in formulating priors on brain activation by simulated data, that serves as a proxy for real-world conditions, possibly capturing a wide range of brain activity patterns[81]. However, these approaches typically incorporate the forward model information only indirectly through their training data, which can limit their accuracy and generalizability in different contexts [68].
@@
To combine worlds, we propose a novel hybrid approach, named 3D-PIUNet, that combines the strengths of both classical methods, which provide physics-informed solutions that generalize across different forward models, and deep learning methods, which learn complex, data-driven priors to improve accuracy. We incorporate physics-informed prior knowledge from the forward model via the pseudo-inverse solution provided by eLORETA [55]. Our method leverages a 3D convolutional neural network to refine this pseudo-inverse solution according to a learned data-driven prior that is able to capture complex patterns in the source space. We show a comparison of classical approaches, end-to-end learning and our proposed 3D-PIUNet in Fig. 1.
@@
Our experimental results on a comprehensive synthetic dataset demonstrate that 3D-PIUNet consistently outperforms both eLORETA and end-to-end deep learning methods in terms of source localization accuracy without requiring temporal information. The network effectively learns to refine the pseudo-inverse solution, leading to more accurate and robust reconstructions, especially in the presence of noise. This work highlights the potential of best-of-both-worlds approaches to overcome the limitations of traditional and data-driven end-to-end methods in EEG source imaging.
@@
  • extensively benchmark our approach in diverse settings, demonstrating its superiority over both classical methods and end-to-end approaches across different signal-to-noise levels, number, and size of sources.
@@ II-A Problem Setup
where a known lead field matrix $\textbf{L}\in\mathbb{R}^{M\times 3N}$ maps the electrical activity of the brain, $\textbf{x}\in\mathbb{R}^{3N\times T}$, to the sensor measurements, $\textbf{y}\in\mathbb{R}^{M\times T}$. Here, $M$ describes the number of sensors on the scalp, $N$ the number of potential brain sources within the brain, with $3$ dimensions to allow free orientation, and $T$ the number of time steps. The noise vector $\boldsymbol{\epsilon}$ describes the measurement noise, covering inaccuracies in the forward model, electrical pollution and sensor errors. This work focuses on the spatial aspect of source localization and considers a single time point, i.e., $T=1$.
@@ II-B Classical Approaches to EEG Source Localization
On the other hand, distributed methods assume that brain activity can be spread across regions involving multiple dipoles. These methods place a dense set of potential source locations within the head volume. The inverse problem becomes a regularized regression problem, predicting the amplitudes of each location, thus not constraining the number of dipoles [49]. The Minimum Norm Estimates (MNE) [28] predicts the distribution of sources by minimizing the $\ell_{2}$-norm of the source amplitudes. The $\ell_{2}$-norm offers a closed-form solution and is, therefore, fast to compute. However, it has a bias towards sources close to the sensors. To counteract this bias, different weighted minimum norm solutions [46] have been proposed, such as dSPM [11], LORETA [54], sLORETA [56] and eLORETA [55] or combinations of those [57]. As this weighting promotes a spatial smoothness prior, small sources might be missed or blurred. To obtain sparse solution with minimal non-zero sources, one can replace the $\ell_{2}$-norm with $\ell_{1}$-norm, called Least Absolute Shrinkage and Selection Operator (LASSO) [72] or selective minimum norm [48]. By combination of sparse and dense prior, it is also possible to generate solutions with mixed spatial extent [32].
@@ II-C Deep Learning Methods
In recent years, learning-based approaches have become popular in the sciences due to their flexibility and fast inference time, including EEG data analysis [60]. A better understanding of brain activity due to virtual brain models [63] and large-scale dynamical models [3] allow the generation of realistic simulation data acting as a powerful data-prior. To learn this data prior, previous work focused on end-to-end learning with different network architectures. For single timestep predictions, fully connected networks have been used to predict source locations directly [53, 37] showing the same limitation in fixed number of sources as dipole fitting. Convolutional networks have been proposed to encode spatial correlation of sensor measurement [35], mixing EEG and MEG signals [41] or to incorporate the time dimension [71]. Other recent approaches use LSTMs for temporal predictions after encoding the measurements either with residual networks [67], graph convolutional networks [40], dilated convolution [80] or multi-head attention [81]. However, all these end-to-end methods use the forward model only during data generation and require retraining or fine-tuning when the forward model changes [68]. Physics-informed machine learning bridges the gap and has shown superior performances in fields like fluid dynamics [58], quantum chemistry [7], quantum circuits [52], and other inverse problems like Computed Tomography (CT) [42] and MRI [38]. Dinh et al. [15] uses an LSTM to refine the pseudo-inverse solution by including temporal context, making a first step towards physic-inspired source localization. However, their training target was restricted to the pseudo-inverse solution, not utilizing the benefits of learning with simulation. In the following, we explain our hybrid approach 3D-PIUNet for source localization, combining the benefits of classical and deep learning methods.
@@ III-A Starting from Pseudo-Inverse Solution
Previous works using neural networks for source estimation [35, 67, 81] learn a mapping $\hat{\textbf{x}}=g_{\theta}(\textbf{y}):\mathbb{R}^{M}\to\mathbb{R}^{3N}$ from sensor to source space. This data-driven approach allows the model to learn complex, non-linear relationships directly from simulated data, potentially capturing patterns that may be difficult to model explicitly using traditional techniques. Even though this end-to-end learning is common practice in modern deep learning, it has some drawbacks for inverse problems. Most importantly, the model needs to learn the physical dynamics only from data, while not having access to a forward model. Therefore, the learned model heavily depends on the lead fields used during training and cannot generalize well to other forward models. For example, when the number of sensors changes, classical end-to-end networks require the training of a new input layer.
We tackle these shortcomings by starting from a pseudo-inverse solution that maps the measurements to the source space. This way, the network can leverage the established physical knowledge embedded in the forward model. Importantly, as long as the pseudo-inverse matrix maintains full column rank, which inverse methods like eLORETA achieve through regularization, this transformation preserves all information contained in the measurement vector. In cases where this condition is not met, some degradation in the reconstruction quality can occur, as the transformation loses some measurement information. During training, the model learns a data prior in source space and can focus on refining the pseudo-inverse solution according to this prior. While this framework has been applied to other inverse problems like CT [42] and MRI [38] reconstruction, to the best of our knowledge, it is the first application to EEG data, while previous work only focused on temporal but not spatial refinement [15]. Given the pseudo-inverse $\textbf{L}^{\dagger}$}, the neural network is trained using the transformed measurements $\tilde{\textbf{x}}={\textbf{L}^{\dagger}}\textbf{y}$ as an input. It then learns the refinement task as a mapping within source space $f_{\theta}:\mathbb{R}^{3N}\to\mathbb{R}^{3N}$:
@@
One can interpret the pseudo-inverse solution as a non-learnable physics-informed input layer that precedes the deep neural network. This way, the pseudo-inverse can be adapted to account for different subjects at inference time, when a subject specific forward model is available. By mapping the measurements into the source space, the input dimensionality does not depend on the sensing system and the same architecture can be used, enabling transfer learning.
@@
Figure 2: An overview of the 3D-PIUNet architecture. First, the eLORETA pseudo inverse $\textbf{L}^{\dagger}$} maps the EEG measurement to a 3D source space of $32\times 32\times 32$ ($32^{3}$) resolution. Next, a series of 3D convolutional residual blocks encodes the pseudo-inverse down to an $8\times 8\times 8$ latent representation. The skip connections between the encoder and decoder allow the model to preserve high-resolution information. At the latent resolution, spatial attention blocks facilitate enhanced interactions between voxels.
@@
We divide the brain sources to a regular 3D grid. This allows the use of a 3D convolutional neural network (CNN) [45]. CNNs excel in this domain due to their parameter efficiency and ability to capture spatial relationships. For tasks like source localization, where preserving spatial information across scales is crucial, the U-Net architecture [59] is particularly well-suited. It employs skip connections between the encoding (downscaling) and decoding (upscaling) blocks, allowing information from the input to propagate directly to the decoder and aiding in the reconstruction of fine details. For this work, we adopt a state-of-the-art U-Net architecture[51]. We illustrate the network architecture of 3D-PIUNet in Fig. 2.
@@
Before applying the network to our pseudo sources $\tilde{x}$, we use zero padding, extending the potential source locations to a regular $32\times 32\times 32$ grid. The architecture begins by lifting the pseudo inverse to a $C=32$ channel feature map, effectively creating multiple representations of the initial source estimate. Next, the network progressively down-samples, by factors of 2, to a bottleneck resolution of ($8\times 8\times 8$) while doubling the feature channels after each down convolution, increasing the receptive field of each convolutional layer. At each resolution, we use two 3D convolutional residual blocks (depicted in blue). Each residual block starts with a Group Norm layer [79], a normalization technique that is less dependent on batch size, leading to improved training stability and performance. This is followed by SiLU activation, a smooth, non-linear activation function outperforming ReLU for deep neural networks [18]. The core computation is done by the $3\times 3\times 3$ convolution layers that aggregate neighboring information to learn descriptive features and refine the source estimates. Each block uses residual connections, where the input of each block is added to its output. This improves the gradient flow during backpropagation and simplifies the optimization problem [34].
After the convolutional blocks reduce the spatial dimensionality, we employ Spatial Attention Blocks (depicted in red) to enhance feature representation. These blocks use three separate linear transformations via $1\times 1\times 1$ convolutions that share weights across all spatial locations, serving as position-wise linear projections. The resulting tensors are then flattened along the spatial dimensions, resulting in $8^{3}=512$ tokens and producing the query $Q\in\mathbb{R}^{512\times 4C}$, key $K\in\mathbb{R}^{512\times 4C}$, and value $V\in\mathbb{R}^{512\times 4C}$ matrices required for the attention mechanism [74]. The attention operation involves computing the matrix product of the query and key, scaled by the square root of the number of channels of query, key and value $d_{k}=4C$, and applying a softmax function to generate the attention map. This map is used to weight the value matrix, thereby enabling each voxel to prioritize the most relevant information from other locations:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$ (3)
@@
The resulting matrix is reshaped back into the 3D volume and is subsequently processed through a $1\times 1\times 1$ convolution for a final linear projection. The resulting feature maps are then added to the input of the attention block, effectively integrating contextual information from across the spatial domain. In contrast to Vision Transformers [16] that represent a group of pixels as a single token, we use a token for each voxel, but for computational reasons only in low-resolution embedding space. The attention mechanism allows the model to focus on the most relevant regions for source reconstruction. These lower layers of the network capture a broader spatial context, and attention allows the model to weigh the importance of different regions within this context, leading to more accurate refinement of the source estimates.
@@ III-C Training Details
[…]
@@
We used the Adam optimizer [44] with a learning rate of 0.00001 and a batch size of 32. The training process consisted of 60000 steps, where the data is generated via the simulation explained in section IV-B. The code is available under https://github.com/MarcoMorik/3D-PIUNet.
@@ IV-A Forward Model
To compute the forward model, we employ a three-shell Boundary Element Method (BEM) [21] approach implemented within the MNE-Python library [24]. To represent potential sources of brain activity, a volumetric source space is constructed. This source space consists of a three-dimensional grid with a uniform voxel spacing of 7 mm. For training, the fsaverage [19] head geometry is used with three concentric layers representing the scalp, skull, and brain tissues and standard conductivity values assigned to these layers: $0.3S/m$ (scalp), $0.006S/m$ (skull), and $0.3S/m$ (brain). These values are based on established studies that provide realistic approximations for human head tissues [35].
To calculate the forward solution, the minimum distance between potential sources and the inner skull surface is set to 5 mm. This parameter helps exclude source locations that are unrealistically close to the skull boundary, resulting in $N=4819$ sources. The volumetric source space is constructed with freely oriented sources, allowing each voxel to represent neural activity with a three-dimensional vector for direction and amplitude, providing greater flexibility in modeling complex neural activity patterns. Sensor positions are defined using a standard Easycap-M10 montage with $M=61$ sensors. The resulting lead field matrix $\textbf{L}\in\mathbb{R}^{(61\times(3\cdot 4819))}$ describes the mapping from source locations and orientation to sensors.
As using the same model for both forward and inverse solutions can lead to over-optimistic results — the so-called inverse crime [9, 43] — we generate a second BEM solution for testing. To account for variations in head geometry, we use a different BEM model generated from MRI data provided by the Open Access Series of Imaging Studies 1 (OASIS1)[47]. For test data generation, the conductivity is changed to $0.332S/m$ (scalp), $0.0113S/m$ (skull), and $0.332S/m$ (brain) following an established setup [35, 78]. Especially the doubling of skull conductivities has a strong influence on the resulting forward model [75]. This ensures that our models are validated against a different forward model, simulating more realistic conditions and enhancing the robustness of our validation process.